OpenStudy (lgg23):

Consider the sequence a(sub n) = n cos(npi)/(2n - 1). Write the first five terms of a(sub n), and find the limit of a(sub n) as n approaches infinity.

OpenStudy (anonymous):

replace n by 1,2,3,4,5 in \[\frac{n\cos(n\pi)}{2n-1}\]

OpenStudy (anonymous):

\(n=1\) get \(\frac{1\cos(1\pi)}{2-1}=-1\)

OpenStudy (anonymous):

\(n=2\) get \(\frac{\cos(2\pi)}{3}=\frac{1}{3}\) and so on